1. Find the area under the t-distribution between from t = - 2.228 to t = 2.228 when n = 11.

A. 0.95

B. 0.995

C. 0.975

D. 0.90


1
Expert's answer
2021-06-22T10:36:06-0400

t = - 2.228 to 2.228

n - 1 = 10

From t-table

Area = 1 - 2t(2.228) = 1 - 2(0.025) = 1 - 0.05 =0.95

Answer: A. 0.95.